A sphere with radius R is uniformly charged with a total charge of Q. The magnitude of the electric field at a distance R/2 from the sphere's centre is:
1
Expert's answer
2020-08-14T16:01:29-0400

The electric field inside sphere:


"E=\frac{kq}{R^3}r"

"E=\frac{kq}{R^3}0.5R=\frac{kq}{2R^2}"